OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−2, 4) and a directrix of y = 6. Put the equation in standard form.

OpenStudy (campbell_st):

ok... so the parabola is concave down since the directrix is above the focus... the distance between the focus and dirextrix is 2 units.... half this distance is the focal length.... so the focal length a = 1 the vertex is at (-2, 4 + a) so the vertex is at (-2, 5) so now using a standard for or the parabola for concave down \[(x - h)^2 = -4a(y - k)\] which has a vertex at (h, k) and a is the focal length, you can say \[(x - (-2))^2 = 4 \times 1 \times(y - 5)\] just simplify for the equation.... hope it helps.

OpenStudy (campbell_st):

the alternate method is to use the distance formula  pick a point on the parabola call it P(x, y) the perpendicular from P to the directrix has coordinates (x, 6) so then the distance from (x, 6) to P(x, y) has to equal the distance from P(x, y) to the focus (-2, 4) so using the distance formulas, you get \[\sqrt{(x - x)^2 + (y - 6)^2} = \sqrt{(x - (-2))^2 + (y - 4)^2}\] sosimplify this to find the equation... hope it helps